# columns=colsep=4pt, rows=rowsep=.5pt

### v1

@@ 4.5 Linear Classification on ImageNet-1k
We obtain strong k-NN performance with our proposed decoupling approach built on top of CARP, demonstrating the benefit of increased prototype spread with a $1.8$-percentage-point improvement in k-NN performance (Table 3). As is well known (Caron et al., 2021), linear evaluation results are highly sensitive to hyperparameter choices, particularly for methods with diverse prototypes (Darcet et al., 2025; Oquab et al., 2024). Consequently, rather than performing extensive hyperparameter tuning for linear classification, we follow a light grid-search protocol [^7] See Appendix C.1 for an extended discussion of the linear evaluation protocol and additional results. and place primary emphasis on the k-NN metric, which provides a more robust, fine-tuning-free assessment of representation quality.
@@ 5 Related Work
Self-Supervised Learning leverages unlabeled data by enabling the model to generate its own supervisory signals. Early approaches introduced pretext tasks such as image inpainting (Pathak et al., 2016) and solving jigsaw puzzles (Noroozi and Favaro, 2016). Since then, a dominant paradigm has emerged that employs joint-embedding architectures within a contrastive framework (Chen et al., 2020, 2021; Chen and He, 2021; Grill et al., 2020), which was later extended by incorporating a predictor network (JEPA) (Assran et al., 2023b). Another line of work departs from joint-embedding architectures, using reconstruction-based objectives such as MAE (He et al., 2022) and BEiT (Bao et al., 2022) based on MIM objectives.
Prototypical Self-Supervised Learning extends joint-embedding methods by introducing a discrete set of learnable prototypes that serve as targets for representation assignment. Such formulations have achieved state-of-the-art performance among computer-vision–only SSL approaches (Siméoni et al., 2025; Venkataramanan et al., 2025), rivaling the effectiveness of language-supervised alternatives (Fan et al., 2025; Radford et al., 2021). Early methods performed cluster assignments once per epoch (Caron et al., 2018; Li et al., 2021), which could result in outdated prototypes. To mitigate this issue, online methods were introduced (Asano et al., 2019; Caron et al., 2020), enabling the joint optimization of the encoder and prototypes, a strategy that has since become the status quo for most prototypical formulations (Assran et al., 2022; Oquab et al., 2024; Siméoni et al., 2025; Silva and Ramírez Rivera, 2022; Caron et al., 2021; Zhou et al., 2022; Assran et al., 2023a; Ruan et al., 2023). However, this joint optimization also introduced the phenomenon of partial prototype collapse which Govindarajan et al. (2023) was first to identify, and has since been addressed to varying degrees of success both explicitly (Govindarajan et al., 2024; Wen et al., 2024) and implicitly (Silva and Ramírez Rivera, 2023; Darcet et al., 2025). Recent work departs from learning prototypes explicitly, instead using fixed high-dimensional codes sampled from a Rademacher distribution (Sansone et al., 2025) or non-parametric targets built from a queue of past encoder representations that act as effective prototypes (Gidaris et al., 2021, 2024; Silva et al., 2024, 2025).
@@ 6 Conclusion
Across extensive experiments, including imbalanced datasets, our decoupling strategy consistently enhances prototype spread, strengthens learned representations, and improves robustness to challenging data distributions. Together, these findings highlight the importance of breaking the joint optimization paradigm and position decoupled prototype learning as a simple, principled, and scalable path forward for prototypical SSL.
@@ Appendix A Online Gaussian Mixture Model
We use a variation of split-and-merge (Ueda and Nakano, 2000) for regularization. But instead of merging low-weight clusters, we instead we apply a method we call split-resurrect. Whenever some component $k$ exceeds a weight threshold, we identify the lightest component $j$, reinitialize its mean (scale-aware random initialization) and reset its variance, and then split the mass of $k$ evenly, $\pi_{k}\leftarrow\pi_{j}\leftarrow\tfrac{1}{2}\pi_{k}^{\text{old}}$. For high cluster weights, we regularize the scale of the dominant mean by
@@ A.1 Ablation: Importance of the Different Components
Responsibility Based Forgetting
(Celaya and Agostini, 2015)
Annealing
(Ueda and Nakano, 1998)
Resurrect
Rescaling
Lin.
@@
72.25
@@ B.1 Hyperparameters
We report the hyperparameters used to train the vision transformer backbones (Dosovitskiy et al., 2021) with CARP and CARP + Decoupling in Table B.1. In addition to these hyperparameters, CARP + Decoupling employs a weight threshold of $0.3$ for the resurrect strategy, which is relatively high given the $65536$ components in the mixture. Furthermore, for the forgetting factor $\eta$, we use a linear scheduler that increases from $0.1$ at the start of training to $0.5$ at its conclusion.
@@ B.1.3 Experiment: Effect of prototype diversity across head, medium and tail classes
We train all our baselines and CARP + Decoupling on iNaturalist2018 dataset using 10 local crops. The iNaturalist2018 challenge is a large scale species classification challenge that has long been the standard dataset for evaluating long-tailed performance. There are 8142 classes in this dataset, with 437,513 training images, and 24,426 validation images. We use the validation set for evaluation as the real test is deliberately left unavailable for public use. The dataset follows a heavy long-tail, with multiple classes appearing very few times in the training set. Following Liu et al. (2019), we consider all classes with more than 100 images each as head classes, between 20 to a 100 images each as medium classes, and less than 20 images each to be tail classes. Using this design choice, we have 842 head classes, 3701 medium classes, and 3599 tail classes.
@@
Appendix C Linear Evaluation
C.1 ImageNet-1k

### v2

@@ 4.5 Linear Classification on ImageNet-1k
We obtain strong k-NN performance with our proposed decoupling approach built on top of CARP, demonstrating the benefit of increased prototype spread with a $1.8$-percentage-point improvement in k-NN performance (Table 3). As is well known (Caron et al., 2021), linear evaluation results are highly sensitive to hyperparameter choices, particularly for methods with diverse prototypes (Darcet et al., 2025; Oquab et al., 2024). Consequently, rather than performing extensive hyperparameter tuning for linear classification, we follow a light grid-search protocol [^6] See Appendix C.2.1 for an extended discussion of the linear evaluation protocol and additional results. and place primary emphasis on the k-NN metric, which provides a more robust, fine-tuning-free assessment of representation quality.
4.6 Transfer Learning
The transfer learning results in Table 4 show that adding decoupling to CARP does not lead to a substantial increase in transfer learning performance. Across ResNet-50 and ViT-Small/16 backbones, the decoupled variants achieve comparable accuracy to the standard CARP, with small variations across datasets. This is consistent with prior observations by Govindarajan et al. (2024), who report that increasing prototype diversity does not necessarily lead to improved transfer performance.
Table 4: Transfer learning results. We follow Ericsson et al. (2021) transfer learning protocol. Top performing in bold, top-2 underlined.
Method Epochs AircrAircr C101 Cars Flower Food Pets SUN397 VOC2007 Avg
ResNet-50
DINO 400 59.95 90.91 65.92 95.63 78.40 89.04 66.08 84.32 78.78
CARP 400 61.03 91.66 64.21 95.82 78.88 90.25 66.10 84.51 79.06
CARP + Decoupling 400 $58.50$ 92.07 67.28 95.91 78.74 91.11 65.24 84.55 79.18
ViT-Small/16
CARP 300 $59.16$ 93.61 64.48 96.40 77.51 93.38 65.94 85.18 79.46
CARP + KP 300 60.48 93.54 63.55 96.44 78.61 93.40 65.98 85.34 79.67
CARP + Decoupling 300 61.78 92.91 63.90 96.20 78.49 94.13 64.53 85.35 79.66
4.7 Memory Usage and Time Consumption
Table 5: Peak memory reserved per GPU for CARP and CARP + Dec. across increasing global batch sizes.
Batch Size CARP CARP + Dec.
128 5.4G 5.3G $\downarrow$ 0.1
256 9.3G 9.1G $\downarrow$ 0.2
512 17.3G 16.8G $\downarrow$ 0.5
1024 32.3G 31.5G $\downarrow$ 0.8
2048 62.4G 60.9G $\downarrow$ 1.5
To assess the memory implications of the proposed decoupling strategy, we measure peak GPU memory reserved per iteration when training a ViT-S/16 model on ImageNet-1k with 10 local crops, comparing the joint-optimized CARP baseline with our decoupled variant while varying the global batch size. All experiments are conducted on AMD Instinct MI250X GPUs. Table 5 reports the per-GPU memory reserved for each approach. At smaller batch sizes (e.g., 128), the reduction in memory footprint is minimal, as activation and optimizer-state requirements dominate overall usage in this regime. However, as the batch size increases, the decoupled formulation yields progressively larger savings. This behavior follows from the mechanics of the online GMM update: because prototype parameters are updated immediately after the forward pass and independently of the encoder’s loss, the method avoids storing the computation graph and intermediate tensors associated with prototype optimization. In contrast, joint-optimization approaches must retain these tensors until backpropagation, contributing to the observed increase in memory usage with larger batch sizes. Because patch-level objectives introduce substantially higher memory requirements, we expect the memory savings from decoupling to be even greater in methods such as iBOT or DINOv2, which we leave as an area for future investigation.
In terms of training time, Table 6 shows that the proposed decoupling method adds only a minimal increase, just 0.2 hours, indicating that the memory savings are achieved without any significant impact on runtime.
Table 6: Training time comparison showing that introducing the online Gaussian Mixture Model for decoupling incurs only a negligible increase of 0.2h.
Method Epochs Batch Size Total Crops Time (h)
CARP 100 1024 12 37.9
CARP + Decoupling 100 1024 12 38.1
@@ 5 Related Work
Self-Supervised Learning leverages unlabeled data by enabling the model to generate its own supervisory signals. Early approaches introduced pretext tasks such as image inpainting (Pathak et al., 2016) and solving jigsaw puzzles (Noroozi and Favaro, 2016). Since then, a dominant paradigm has emerged that employs joint-embedding architectures within a contrastive framework (Chen et al., 2020; 2021; Chen and He, 2021; Grill et al., 2020), which was later extended by incorporating a predictor network (JEPA) (Assran et al., 2023a). Another line of work employs reconstruction-based objectives instead (He et al., 2022; Bao et al., 2022).
Prototypical Self-Supervised Learning extends joint-embedding methods by introducing a discrete set of learnable prototypes that serve as targets for representation assignment. Such formulations have achieved state-of-the-art performance among vision-only SSL approaches (Siméoni et al., 2025; Venkataramanan et al., 2025), rivaling the effectiveness of language-supervised alternatives (Fan et al., 2025; Radford et al., 2021). Early methods performed cluster assignments once per epoch (Caron et al., 2018; Li et al., 2021), which could result in outdated prototypes. To mitigate this issue, online methods were introduced (Asano et al., 2020; Caron et al., 2020), enabling the joint optimization of the encoder and prototypes, a strategy that has since become the status quo for most prototypical formulations (Assran et al., 2022; Oquab et al., 2024; Siméoni et al., 2025; Silva and Ramírez Rivera, 2022; Caron et al., 2021; Zhou et al., 2022; Assran et al., 2023b; Ruan et al., 2023). However, this joint optimization also introduced the phenomenon of partial prototype collapse which Govindarajan et al. (2023) was first to identify, and has since been addressed to varying degrees of success both explicitly (Govindarajan et al., 2024; Wen et al., 2024) and implicitly (Silva and Ramírez Rivera, 2023; Darcet et al., 2025). Recent work departs from learning prototypes explicitly, instead using fixed high-dimensional codes sampled from a Rademacher distribution (Sansone et al., 2025) or non-parametric targets built from a queue of past encoder representations that act as effective prototypes (Gidaris et al., 2021; 2024; Silva et al., 2024; 2025).
Limitations: Although our proposed decoupling approach alleviates prototype collapse, it introduces several limitations. The current implementation lacks a mechanism to ensure that prototype updates track the encoder’s rate of change, as updates rely solely on the forgetting factor $\eta$, which cannot adapt to shifts in the online estimation of the mixture. The use of an online GMM also adds hyperparameters, and it assumes that incoming representations follow a Gaussian distribution. While our experiments show that the method eliminates prototype collapse across all investigated thresholds, the complete removal of collapse is not universally beneficial across frameworks (cf. Appendix A.3).
@@ 6 Conclusion
Across extensive experiments, including imbalanced datasets, our decoupling strategy consistently enhances prototype spread, while its impact on downstream representations and robustness to challenging data distributions varies across architectures. Together, these findings highlight the benefits of breaking the joint optimization paradigm, demonstrating that decoupled prototype learning is a simple, principled, and scalable strategy for preventing collapse and improving representation quality in several prototypical SSL settings.
Acknowledgments
This work was funded, in part, by RCN (the Research Council of Norway) through Visual Intelligence, Centre for Research-based Innovation (grant no. 309439), and FRIPRO (grants no. 315029 and 359216). The computations were performed, in part, on resources provided by Sigma2—the National Infrastructure for High-Performance Computing and Data Storage in Norway (Project NN8104K). We thank Shashanka Venkataramanan and Yuki M. Asano (Franca team) for providing their reproduced DINOv2 prototype weights, which broadened the scope of our study.
@@ Appendix A Online Gaussian Mixture Model
Throughout training, we use a variation of split-and-merge (Ueda and Nakano, 2000) for regularization. But instead of merging low-weight clusters, we instead we apply a method we call split-resurrect. Whenever some component $k$ exceeds a weight threshold, we identify the lightest component $j$, reinitialize its mean (scale-aware random initialization) and reset its variance, and then split the mass of $k$ evenly, $\pi_{k}\leftarrow\pi_{j}\leftarrow\tfrac{1}{2}\pi_{k}^{\text{old}}$. For high cluster weights, we regularize the scale of the dominant mean by
@@ A.1 Ablation: Importance of the Different Components
Responsibility Based Forgetting (Celaya and Agostini, 2015) Annealing (Ueda and Nakano, 1998) Resurrect Rescaling Lin.
@@
[…]
A.2 Ablation: Sensitivity Analysis of the Forgetting-Factor Scheduler
The forgetting factor $\eta$ used in the online GMM update is defined through a scheduler of the form
$\eta=0.99-\frac{1}{at-b},$ (A.7)
where $t$ denotes the training timestep, $a$ controls the rate at which ($\eta$) approaches its upper bound of $0.99$, and $b=\frac{1}{0.99-c}$. The parameter $c$ determines the initial value of the scheduler, while $a$ governs its progression over the course of training.
In the main experiments, we set $a=0.001$ and $c=0.89$, which yields an initial forgetting factor of $\eta=0.89$ at the first training step. To assess the sensitivity of this choice, we evaluated alternative schedules with different values of $a$ and $c$ on ImageNet-1k using a ViT-S trained for 100 epochs with 6 local crops, varying each parameter independently while keeping all other hyperparameters fixed.
Table A.2: Ablation of forgetting-factor scheduler parameters on ImageNet-1k with CARP+Dec..
Method Arch Epoch $a$ $c$ k-NN (%)
CARP + Dec. ViT-S/16 100 0.001 0.75 66.82
CARP + Dec. ViT-S/16 100 0.001 0.95 67.41
CARP + Dec. ViT-S/16 100 0.0005 0.89 69.26
CARP + Dec. ViT-S/16 100 0.002 0.89 68.99
CARP + Dec. ViT-S/16 100 0.001 0.89 69.52
Overall, the results indicate that performance remains stable across a reasonable range of scheduler configurations. Adjusting the intercept $c$ produces the largest effect: decreasing $c$ to $0.75$ leads to a notable drop in accuracy, as the resulting schedule causes the prototypes to update too aggressively early in training. Increasing $c$ to $0.95$ slows the early updates and yields a milder degradation, but still performs worse than the baseline. The intermediate value $c=0.89$ provides a balanced update rate that appears most effective. In contrast, modifying the convergence rate through $a$ has a more negligible influence on performance, suggesting that the schedule is less sensitive to this parameter.
A.3 Limitations
Although our proposed decoupling approach alleviates prototypical collapse, it comes with certain limitations. First, the current implementation lacks a mechanism to ensure that prototype updates remain aligned with the encoder’s rate of change. Prototype updates are governed solely by the forgetting factor $\eta$, and although a linear schedule performs reasonably well, it does not allow the method to adjust adaptively to the online estimation of the Gaussian mixture. Because $\eta$ plays a role analogous to a learning rate in gradient-based optimization, an interesting direction for future work would be to incorporate momentum to better regulate prototype updates.
Furthermore, incorporating an online GMM to enable our decoupling strategy introduces additional hyperparameters, such as the smoothing factor $\beta$ and the forgetting factor $\eta$, which are required to manage the sparse update regime associated with the large prototype sets used in prototypical SSL frameworks. The online mixture model also assumes that incoming representations follow a Gaussian distribution with meaningful variation in magnitude. This assumption aligns with CARP (Silva and Ramírez Rivera, 2023), which does not normalize its projection head outputs, but it does not strictly hold for DINO (Caron et al., 2021), where the outputs are normalized. Although our internal experiments did not reveal noticeable differences when replacing the Gaussian with a von Mises-Fisher distribution, the current formulation of our decoupling method is built around a Gaussian assumption.
Lastly, our experiments in Section 4.2 and Section 4.3 show that the proposed decoupling method effectively eliminates prototype collapse across all our investigated thresholds $\epsilon$ following Definition 2.1. However, the absence of prototype collapse is not universally beneficial for all frameworks, as evidenced by the results in Section 4.4. We nonetheless view this as a meaningful contribution to the community, as it enables systematic evaluation of prototypical SSL methods across different degrees of collapse: whereas Govindarajan et al. (2024) reduces collapse, our approach eliminates it entirely [^7] Under all thresholds used to define unique prototypes in our evaluations., providing a complementary perspective for understanding its role in representation learning.
@@ B.1 Hyperparameters
We report the hyperparameters used to train the vision transformer backbones (Dosovitskiy et al., 2021) with CARP and CARP + Decoupling in Table B.1. In addition to these hyperparameters, CARP + Decoupling employs a weight threshold of $0.3$ for the resurrect strategy, which is relatively high given the $65536$ components in the mixture. Furthermore, for the smoothing factor $\beta$, we use a linear scheduler that increases from $0.1$ at the start of training to $0.5$ at its conclusion.
@@ B.1.3 Experiment: Effect of prototype diversity across head, medium and tail classes
We train all our baselines and CARP + Decoupling on iNaturalist2018 dataset using 10 local crops. All models use a ViT-S/16 backbone trained for 300 epochs on iNaturalist2018. The iNaturalist2018 challenge is a large scale species classification challenge that has long been the standard dataset for evaluating long-tailed performance. There are 8142 classes in this dataset, with 437,513 training images, and 24,426 validation images. We use the validation set for evaluation as the real test is deliberately left unavailable for public use. The dataset follows a heavy long-tail, with multiple classes appearing very few times in the training set. Following Liu et al. (2019), we consider all classes with more than 100 images each as head classes, between 20 to a 100 images each as medium classes, and less than 20 images each to be tail classes. Using this design choice, we have 842 head classes, 3701 medium classes, and 3599 tail classes.
@@
Appendix C Extended Results
The preceding sections establish the empirical evidence motivating the proposed decoupled framework and its effects on prototype diversity, training dynamics, and performance across distribution regimes. These findings demonstrate that joint optimization induces a form of shortcut learning, driving prototypes toward redundant representations early in training, whereas decoupled updates maintain a higher prototype diversity.
C.1 Theoretical Perspective on Prototype Collapse
The empirical results consistently reveal a strong link between the way prototypes are optimized and the emergence of partial prototype collapse. These patterns point toward structural properties of the optimization that influence how prototypes distribute across the representation space. This section introduces a theoretical perspective that explains why joint updates promote redundancy and why separating prototype estimation leads to a diverse set of prototypes.
C.1.1 Decoupling and Prototype Collapse
Consider the prototype subproblem at fixed encoder parameters $\theta$ and soft assignments $\gamma_{ik}$ induced by the teacher-student logits. We can pose the subproblem as
$J(C\mid\theta,\gamma)=\sum_{i,k}\gamma_{ik}\|h_{i}-c_{k}\|^{2},\quad h_{i}=f_{\theta}(x_{i}),$
whose unique minimizers are the responsibility-weighted centroids $c_{k}^{*}=\frac{\sum_{i}\gamma_{ik}h_{i}}{\sum_{i}\gamma_{ik}}$. If two distinct components $j\neq k$ have different responsibilities $\gamma_{:j}\neq\gamma_{:k}$, then $c_{j}^{*}\neq c_{k}^{*}$ almost surely; hence collapsed prototypes are not stationary points of the prototype subproblem $J$.
The decoupled optimization approximates a coordinate-descent / EM scheme. For fixed $\theta_{t}$ at step $t$ the GMM update performs an online M-step that pushes the full set of prototypes $C_{t}$ toward the centroids $c_{k}^{*}(\theta_{t},\gamma_{t})$. Moreover, for fixed $C_{t}$, we update $\theta$ on the SSL objective $\mathcal{L}_{f}(\theta,C_{t})$. This keeps prototypes close to block-optimal solutions of the prototype subproblem at each step, which disfavors collapse whenever components have distinct assignments.
In contrast, joint gradient updates on $(\theta,C)$ only enforce joint stationarity of $\mathcal{L}_{f}(\theta,C_{t})$, and do not enforce optimality of $J(C\mid\theta,\gamma)$. In other words, collapse with $c_{j}=c_{k}$ can be joint stationary points of the coupled system, even though they are suboptimal for the prototype subproblem. This is what we observe in our analysis of DINO-style modeling (Table 1, Figure 1a, Figure 4).
We note that if student prototypes of DINO exhibit collapse, the gradients for collapsed components receive identical updates, so collapsed prototypes cannot re-separate. Furthermore, any collapse in the student model will invoke a collapse in the teacher, whose prototypes are EMA updates of the student. Hence, teacher asymmetry does not meaningfully help prevent collapse.
C.1.2 Formalisation of decoupling prototype collapse
Lemma C.1 (Unique Centroids).
For the prototype subproblem
$J(C\mid\theta,\gamma)=\sum_{k}J_{k}(c_{k}),\quad J_{k}(c_{k})=\sum_{i}\gamma_{ik}\|h_{i}-c_{k}\|^{2},$ (C.1)
each $J_{k}$ is strictly convex quadratic in $c_{k}$ with unique minimizer
$c_{k}^{*}(\theta,\gamma)=\frac{1}{m_{k}}\sum_{i}\gamma_{ik}h_{i},\quad m_{k}=\sum_{i}\gamma_{ik}.$ (C.2)
Proof.
The gradient and Hessian of $J_{k}$ are given by
$\nabla_{c_{k}}J_{k}(c_{k})=-2\sum_{i}\gamma_{ik}h_{i}+2m_{k}c_{k},\quad\nabla^{2}_{c_{k}}J_{k}(c_{k})=2m_{k}I_{d}.$ (C.3)
As $m_{k}>0$, the Hessian is positive definite, so $J_{k}$ is strictly convex with unique minimizer
$\nabla_{c_{k}}J_{k}(c_{k})=0\Rightarrow c_{k}^{*}=\frac{1}{m_{k}}\sum_{i}\gamma_{ik}h_{i},$ (C.4)
as we wanted to show. ∎
Proposition C.2 (Non-collapse for decoupled prototype subproblem).
Assume features $h_{i}\in\mathbb{R}^{d}$ are drawn from a distribution that is absolutely continuous w.r.t. Lebesgue measure. Fix responsibilities $\gamma_{ik}\geq 0$ and masses $m_{k}=\sum_{i}\gamma_{ik}>0$.
Then for any pair $j\neq k$ with $\gamma_{:j}\neq\gamma_{:k}$, we have that $\mathrm{Pr}(c_{j}^{*}=c_{k}^{*})=0$, i.e. $c_{k}^{*}\neq c_{j}^{*}$ almost surely. Equivalently, collapsed centroids occur on a subset of measure zero.
Proof.
By Lemma 1, the centroids are
$c_{j}^{*}=\frac{1}{m_{j}}\sum_{i}\gamma_{ij}h_{i},\quad c_{k}^{*}=\frac{1}{m_{k}}\sum_{i}\gamma_{ik}h_{i}.$
These are convex combinations of the features with normalized weight distributions $(\gamma_{ij}/m_{j})_{i}$ and $(\gamma_{ik}/m_{k})_{i}$.
Since $\gamma_{:j}\neq\gamma_{:k}$, these weight distributions differ: there exists at least one index $i^{*}$ such that $\gamma_{i^{*}j}/m_{j}\neq\gamma_{i^{*}k}/m_{k}$.
Collapse $c_{j}^{*}=c_{k}^{*}$ occurs if and only if
$\sum_{i}\frac{\gamma_{ij}}{m_{j}}h_{i}=\sum_{i}\frac{\gamma_{ik}}{m_{k}}h_{i}\quad\Leftrightarrow\quad\sum_{i}w_{i}h_{i}=0,$
where $w_{i}=\gamma_{ij}/m_{j}-\gamma_{ik}/m_{k}$.
Since the weight distributions differ, $w\neq 0$ (i.e., not all $w_{i}=0$). The collapse condition $\sum_{i}w_{i}h_{i}=0$ therefore imposes a non-trivial linear constraint on the features.
Geometrically, this defines a hyperplane in $\mathbb{R}^{n\times d}$; the set $\{(h_{1},\ldots,h_{n}):\sum_{i}w_{i}h_{i}=0\}$ is a (nd1)(nd-1)-dimensional affine subspace of the ambient $nd$-dimensional space. Since each $h_{i}$ is drawn independently from a distribution absolutely continuous with respect to Lebesgue measure, the joint distribution of $(h_{1},\ldots,h_{n})$ is absolutely continuous on $\mathbb{R}^{n\times d}$, and assigns measure zero to any proper affine subspace.
Therefore, $P(c_{j}^{*}=c_{k}^{*})=0$. ∎
Corollary C.3 (No Prototype Collapse at Decoupled Fixed Points).
Let $(\theta^{*},C^{*})$ be a decoupled fixed point, such that
  • $C^{*}$ is a local minimizer of the prototype subproblem
    $C^{*}=\arg\min_{C}J(C\mid\theta^{*},\gamma^{*}),$
  • $\theta^{*}$ is a stationary point of the SSL objective
    $\nabla_{\theta}\mathcal{L}_{f}(\theta^{*},C^{*})=0.$
Then given the conditions of Proposition 1, for any pair of distinct components $j\neq k$ with $\gamma_{:j}^{*}\neq\gamma_{:k}^{*}$, we have
$P(c_{j}^{*}=c_{k}^{*})=0.$
In other words, decoupled fixed points do not admit prototype collapse almost surely whenever the induced responsibilities are non-identical and non-degenerate.
Proof.
Since $C^{*}$ minimizes $J(\cdot\mid\theta^{*},\gamma^{*})$, each centroid must be the responsibility-weighted mean:
$c_{k}^{*}=\frac{1}{m_{k}^{*}}\sum_{i}\gamma_{ik}^{*}h_{i}^{*}.$
By Proposition 1, if $\gamma_{:j}^{*}\neq\gamma_{:k}^{*}$ and $m_{j}^{*},m_{k}^{*}>0$, then
$P(c_{j}^{*}=c_{k}^{*})=0.$
Note that this result does not necessarily hold for coupled optimization such as with DINO, which can yield cases where a joint optima over the SSL loss function admits $c_{j}=c_{k}$.
C.2 Linear Classification
C.2.1 ImageNet-1k